# Zonal vs. Nodal Pricing: An Analysis of Different Pricing Rules in the German Day-Ahead Market

### v1

@@ Abstract
The European electricity market is based on large pricing zones with a uniform day-ahead price. The energy transition leads to shifts in supply and demand and increasing redispatch costs. In an attempt to ensure efficient market clearing and congestion management, the EU Commission has mandated the Bidding Zone Review (BZR) to reevaluate the configuration of European bidding zones. Based on a unique data set published in the context of the BZR, we compare various pricing rules for the German power market. We compare market clearing and pricing for national, zonal, and nodal models, including their generation costs and associated redispatch costs. Moreover, we investigate different non-uniform pricing rules and their economic implications for the German electricity market. Our results indicate that the differences in the average prices in different zones are small. The total costs across different configurations are similar and the reduction of standard deviations in prices is also small based on this data set. A nodal pricing rule leads to the lowest total costs. We also analyze the quality of different pricing rules and their differences with respect to the quality of the price signals and the necessary uplift payments. While the study focuses on Germany, the analysis is relevant beyond and feeds into the broader discussion about pricing rules.
@@
[1]organization=School of Computation, Information and Technology, Technical University of Munich,addressline=Boltzmannstr. 3, city=Garching, postcode=85748, country=Germany
@@ 1 Introduction
A fundamental problem in electricity market design is to compute adequate price signals in the presence of non-convex preferences of market participants. Prices should reflect the true value of electricity and provide the right bidding and investment incentives to all relevant market participants (Cramton,, 2017). In liberalized markets, day-ahead electricity prices are derived from welfare maximization problems, but pricing is challenging due to the non-convexity of these optimization problems (Liberopoulos & Andrianesis,, 2016). A central discussion on electricity market pricing revolves around zonal and nodal pricing (aka. locational marginal pricing). Among the electricity wholesale markets implementing nodal pricing are Argentina, Chile, Mexico, New Zealand, Peru, Russia, Singapore, and several regions in the United States (U.S.). In contrast, a zonal pricing approach has been employed in the large and coupled European markets to reduce the computational burden of market clearing and pricing. Large groups of nodes are aggregated into zones, and the market clearing problem only considers flow constraints between zones. With some exceptions (e.g., Italy, Norway, Sweden), there is a single price zone within a country, and the day-ahead market provides a uniform national electricity price. Ignoring transmission constraints can lead to dispatch decisions that are not physically feasible. Thus, after the market clearing, transmission operators conduct redispatch to ensure that the final allocation aligns with physically feasible power flows.
@@
Due to climate change and the ongoing energy transition, the zonal pricing paradigm has come under scrutiny (Eicke & Schittekatte,, 2022; Bertsch et al.,, 2016; Trepper et al.,, 2015). Volatile renewable energy sources such as wind and solar lead to congestion in different parts of the electricity grid (Neuhoff et al.,, 2013). In addition, the energy transition can lead to structural problems. We will focus on Germany as a case in point. Wind farms have grown significantly in Northern Germany, while much of the industrial demand is in the south. Energy is cleared as if there were no transmission constraints, but the traded energy can often not be delivered. As a result, redispatch costs have increased drastically in recent years to 4.2 billion in 2022 (Bundesnetzagentur,, 2022). The redispatch costs in 2022 were also influenced by the war in the Ukraine, but have been increasing consistently over several years. In an attempt to ensure efficient market clearing and congestion management, the EU Commission – as part of the Clean Energy Package – has mandated a Bidding Zone Review (BZR) process to reevaluate the configuration of European bidding zones. An integral part of this process is a locational marginal pricing (LMP) study conducted by the European Network of Transmission System Operators for Electricity (ENTSO-E) as a basis to identify structural congestion and modified bidding zones. Based on this study, in August 2022, the European Union Agency for the Cooperation of Energy Regulators (ACER) decided on a set of alternative bidding zone configurations based on proposals by ENTSO-E (ACER, 2022a, ). Among others, the BZR considers a split of the German price zone.
In their LMP study, ENTSO-E solved linearized unit commitment models with marginal pricing to obtain nodal prices (ENTSO-E,, 2022). The data is the main input for the Bidding Zone Review and constitutes the results of a multi-year effort of the European TSOs. The discussion of nodal and zonal prices has a long history, but this unique data set allows for a unique opportunity to compare different pricing rules on realistic data and estimate the differences. Note that the allocation considers assets individually at each node and abstracts from block orders or portfolio bids that are common practice in Europe today. The computed prices served as input for clustering approaches to yield alternative bidding zone configurations. Following Article 16 of the BZR methodology, ENTSO-E has published non-confidential data related to the LMP study (ENTSO-E,, 2023). We leverage this dataset to expand on the results of the LMP study, focusing on the German bidding zone. Beyond the choice between four alternative zonal configurations of the German market, our goal is to understand how these configurations compare to nodal prices and the current market with a uniform national price.
Specifically, we compare market clearing and pricing for national, zonal, and nodal models, the generation costs and associated redispatch costs. We consider a redispatch model that minimizes cost-based compensations provided to generators for any changes in their dispatched quantities during redispatch. Moreover, we investigate different non-uniform pricing rules and their economic implications for the German electricity market. We consider Integer Programming (IP) pricing (O’Neill et al.,, 2005) and Convex Hull (CH) pricing (Hogan & Ring,, 2003; Gribik et al.,, 2007) as two established pricing rules in U.S. markets, and the recently suggested Join pricing rule (Ahunbay et al.,, 2022). We also implemented a version of the Euphemia algorithm currently used in European day-ahead markets. In contrast to IP, CH, and Join pricing, Euphemia sacrifices welfare maximization to achieve linear and anonymous prices and budget balance. We compare the outcomes of clearing and pricing rules based on the total costs, average prices, make-whole payments (to compensate paradoxically accepted bids), and lost opportunity costs. For Euphemia, we also compute the welfare loss incurred by the algorithm. We report aggregate results using the generation and demand scenarios for 2025 and climate year data of 2009 provided by the ENTSO-E LMP study. We also focus on individual days to illustrate our findings and their economic implications.
@@ 1.2 Contributions
Our study suggests that a nodal pricing rule lowers system costs as costly redispatch can be avoided. Conversely, by disregarding transmission constraints, the zonal configurations lead to an inefficient dispatch. Even if redispatch is conducted at minimal costs, this inefficiency leads to higher overall system costs. The average price differences across configurations between 1 and 4 zones for Germany are less than 3 EUR/MWh. This could partly be attributed to the capacities assumed by ENTSO-E for 2025, e.g., new solar capacities in the southern region. There are also individual days where the price difference between zones is more than 10 EUR/MWh, but on average the differences are small. The higher total costs for national and zonal prices arise from the fact that due to the omission of network constraints, an inefficient allocation is picked. This results in the selection of generators that would not be part of an efficient nodal dispatch. These inefficiencies lead to higher total costs after redispatch that is required to achieve a feasible dispatch respecting the transmission capacities. As indicated earlier, in the past there was often an oversupply of renewable energy sources in the north and high demand in the south of Germany, which raised expectations that redispatch decreases significantly with two price zones. However, our results based on data from ENTSO-E, does not confirm such expectations.
The redispatch costs in our model do not decrease considerably with a split in 2, 3, or 4 zones. The way how these redispatch costs are computed can be considered as a lower bound. For example, we assume that every generator can be redispatched, but this is not the case in reality. A key goal for the Bidding Zone Review is to find pricing zones such that the variance in prices is small, which is seen as a proxy for network congestion. We show that the differences in the standard deviation of prices caused by congestion in different zones are small compared to the overall standard deviation in prices, which explains why the differences in the redispatch costs are small across the different zonal configurations.
We also provide an analysis of different non-uniform pricing rules such as IP pricing (as used in U.S. power markets), Convex Hull Pricing, and the Join pricing rule. The results illustrate the trade-offs that can be expected in terms of quality of the congestion signals and minimum make-whole payments and provide guidance regarding the analysis of non-uniform pricing rules in the context of CACM 2.0 (All NEMO Committee, 2023b, ). While our analysis focuses on Germany, the policy implications go beyond. The study highlights the challenges in delineating price zones in a zonal market and the trade-offs between different pricing rules based on a new and unique data set that results from a multi-year effort by the European TSOs.
1.3 Limitations and Organization
@@
Like any empirical analysis of its kind, also our study faces limitations. First, our examination focuses on the German day-ahead market without incorporating neighboring countries, cross-border trades, or loop flows through other countries. Second, we do not model the intraday or forward markets, which hold considerable significance within the European electricity system. Third, we assume fixed demand from the ENTSO-E dataset, neglecting the potential for demand reduction incentives under nodal prices. Moreover, the current European market is characterized by portfolio bidding, differing from the unit-commitment bids underlying both the study by ENTSO-E and our analysis. Portfolio bids allow optimization of trades across assets and differ from those feasible in nodal market designs. Incorporating such considerations into our analysis would require making strong and potentially unwarranted assumptions regarding market participant behavior. Finally, the data set provided by ENTSO-E was constructed for the target year 2025, and even though different climate years are considered, it is important to acknowledge that analyses can vary based on differing assumptions regarding supply and demand. However, the ENTSO-E dataset represents the best source available today for such a study. Additional assumptions made for our analysis are described in Section 3. We also emphasize that the discussion surrounding zonal and nodal pricing extends beyond dispatch and prices. In the ongoing Bidding Zone Review, a comprehensive approach is adopted which considers factors such as liquidity, security of supply, transition costs, and investments in low-carbon technologies, which we do not in our analysis.
@@ 2.2 Market Clearing
In practice, market operators employ different bidding languages to encode $v_{b}$ and $c_{s}$, as well as different approximations $\Psi$ of the power flow constraints. In this work, we use the data released for the European bidding zone review (BRZ) to compare allocation and prices under different market clearing mechanisms.
In full generality, the market operator seeks an allocation $(x,y)$, where $x=(x_{b})_{b\in B}$, $x_{b}\in\mathbb{R}^{N\times T}$ is the allocation vector of buyers and $y=(y_{s})_{s\in S}$,$y_{s}\in\mathbb{R}^{N\times T}$ is the allocation vector of sellers. The market operator seeks an allocation that maximizes welfare, by solving the following optimization model:
@@
On European day-ahead markets, buyers and sellers submit hourly bids and block orders (NEMO Committee,, 2019). The BZR data, however, does not contain bids of this structure. Instead, it provides generator and load characteristics similar to unit commitment problems in U.S. markets. Specifically, for buyers $b$ there is only information on a fixed demand profile $P_{b}\in\mathbb{R}^{N\times T}$, while price-elastic bids are unavailable. Consequently, the valuation function of $b$ simplifies to strict demand satisfaction.
@@
$\Psi^{DC}=\{(x,y):\sum_{s\in S}y_{s,nt}-\sum_{b\in B}x_{b,nt}=\sum_{m\in N}-B_{nm}(\theta_{n}-\theta_{m})\;\forall n\in N,t\in T\}$ (4)
In (4), $B_{nm}$ is the susceptance of the line connecting nodes $n$ and $m$. If such a line does not exist, then $B_{nm}=0$, and $B_{nn}=0\;\forall n$. The voltage angle decision variable for node $n$ is denoted as $\theta_{n}$. The constraint set in (4) can be enriched by further constraints, such as thermal line limits, box constraints on the voltage angles, and more (Li et al.,, 2022). We use such a standard DCOPF model in our analysis, similar to what was used in the ENTSO-E study. When the representation $\Psi$ of the transmission network considers all nodes in the transmission network, we refer to nodal market clearing. When only a subset of the grid is considered, we speak of zonal market clearing, as discussed in the following.

### v4

@@ Abstract
The European electricity market is based on large pricing zones with a uniform day-ahead price. The energy transition leads to changes in supply and demand and increasing redispatch costs. In an attempt to ensure efficient market clearing and congestion management, the EU Commission has mandated the Bidding Zone Review (BZR) to reevaluate the configuration of European bidding zones. Based on a unique data set published in the context of the BZR for the target year 2025, we compare various pricing rules for the German power market. We compare market clearing and pricing for different zonal and nodal models, including their generation costs and associated redispatch costs. In numerical experiments with this dataset, the differences in the average prices in different zones are low. Congestion arises as well, but not necessarily on the cross-zonal interconnectors. The total costs across different configurations are similar and the reduction of standard deviations in prices is also small. This might be different with other load and generation scenarios, but the BZR data is important as it was created to make a decision about splits of the existing bidding zones. Nodal pricing rules lead to the lowest total cost. We also evaluate differences of nodal pricing rules with respect to the necessary uplift payments, which is relevant in the context of the current discussion on non-uniform pricing in the EU. While the study focuses on Germany, the analysis is relevant beyond and feeds into the broader discussion about pricing rules in non-convex markets.
keywords:
Nodal Pricing , Zonal Pricing , Non-Convex Markets , Redispatch
journal: Operations Research
@@
[1]organization=School of Computation, Information and Technology, Technical University of Munich,Department and Organization addressline=Boltzmannstr. 3, city=Garching, postcode=85748, country=Germany
@@ 1 Introduction
A central discussion on electricity market pricing revolves around zonal and nodal pricing (aka. locational marginal pricing). Among the electricity wholesale markets implementing nodal pricing are Argentina, Chile, Mexico, New Zealand, Peru, Russia, Singapore, and several regions in the United States (U.S.). In contrast, a zonal pricing approach has been employed in the large and coupled EU market. Large groups of nodes are aggregated into zones, and the market clearing problem only considers flow constraints between zones. It is supposed that no or only limited congestion occurs inside a zone. With some exceptions (e.g., Italy, Norway, Sweden), there is a single price zone within a country, and the day-ahead market provides a uniform national electricity price. Ignoring transmission constraints can lead to dispatch decisions that are not physically feasible. Thus, after the market clearing, transmission operators conduct redispatch to ensure that the final allocation aligns with physically feasible power flows. In the USA, all Independent Service Operators (ISOs) moved from a zonal to a nodal model, partly due to steeply increasing redispatch costs. Aravena et al., (2021) provide an excellent summary of historical developments in different juristictions.
@@
Due to climate change and the ongoing energy transition, the zonal pricing paradigm has further come under scrutiny (Eicke & Schittekatte,, 2022; Bertsch et al.,, 2016; Trepper et al.,, 2015). Volatile renewable energy sources such as wind and solar lead to congestion in different parts of the electricity grid (Neuhoff et al.,, 2013). In addition, the energy transition can lead to structural problems. We will focus on Germany as a case in point. Energy is cleared as if there were no transmission constraints in Germany, but the traded energy can often not be delivered due to transmission constraints. As a result, redispatch volumes and costs have increased substantially in recent years to 34 TWh and EUR 3.1 billion in 2023 (Bundesnetzagentur,, 2024). In an attempt to ensure efficient market clearing and congestion management, the EU Commission – as part of the Clean Energy Package – has mandated a Bidding Zone Review (BZR) to reevaluate the configuration of European bidding zones. An integral part of this process is a locational marginal pricing (LMP) study conducted by the European Network of Transmission System Operators for Electricity (ENTSO-E) as a basis to identify structural congestion and modified bidding zones. Having accurate data and grid models is central for the quality of the BZR. The LMP study constitutes the results of a multi-year effort of the European TSOs to provide a representative data set, and it is unparalleled in terms of the quality and degree of detail of the data. The study is based on an up-to-date model of the electricity grid and supply and demand estimates for 2025 and different climate scenarios. In this study, ENTSO-E solved linearized unit commitment models with marginal pricing to obtain nodal prices (ENTSO-E,, 2022). Based on the resulting prices, in August 2022, the European Union Agency for the Cooperation of Energy Regulators (ACER) decided on a set of alternative bidding zone configurations (ACER, 2022a, ). Following Article 16 of the BZR methodology, ENTSO-E has published non-confidential data related to the LMP study (ENTSO-E,, 2023). We refer to this data set as the BZR data in what follows and leverage this dataset to expand on the results of the LMP study, focusing on the German bidding zone. Germany is particularly interesting, because ACER suggested the evaluation of several configurations with splits of the large German-Luxembourgish bidding zone.
1.2 Related Literature
A few recent studies aimed to analyze the impact of a bidding zone split on Germany or other countries in Europe. These studies are typically based on proprietary data sets and models and they assume different target years. The consulting company Aurora published results of another study where they predict 5 Euros per MWh price difference between the north and the south of Germany in the years 2030, increasing to 9 Euros in 2045.111https://auroraer.com/insight/power-market-impact-of-splitting-the-german-bidding-zone/ Models and data are proprietary. In another study by the EWI and THEMA consulting [^2] https://www.ewi.uni-koeln.de/en/publications/price-impact-of-a-german-bidding-zone-split/ the authors predict differences in prices in a southern and a northern zone of around 15 to 10 Euros per MWh between 2024 and 2033, with a decreasing trend. The study is based on two separate models and data by the EWI and THEMA consulting that are not publicly available. The size of the spread differs in THEMA’s and EWI’s results according to the authors. For 2030 the study makes a number of assumptions regarding electrolysers and new lines being built, batteries, and a shift of renewables from North to South. Tiedemann et al., (2024) discuss the market value of renewable energy sources after a zonal split for 2030. Their study is based on a data set and grid model by the Fraunhofer IEE and they predict around 10 Euros per MWh difference between the north and south of Germany respectively in 2030. They do not comment on reductions in redispatch by a zonal split.
It is worth mentioning that the data sets used in these studies and the underlying grid models differ substantially, and so do the target years analyzed. For example, there are differences in the number of lines in the BZR data set compared to the widely used JAO Static Grid model. [^3] https://www.jao.eu/static-grid-model,accessedinMay2024. Also, the assumptions made for supply and demand for the future differ. We stick to the BZR dataset which appears to have a very comprehensive grid model and a near-term target year of 2025.444There are also earlier studies on a zonal split, such as those by Trepper et al., (2015), but the data sets and assumptions are even harder to compare due to the significant changes in the past ten years.
All of these studies analyze static models where demand and supply are given as we do. Ambrosius et al., (2020) use a multi-level mixed-integer non-linear model based on Grimm et al., (2019) to analyze investment incentives over time. Such models cannot accommodate the level of detail in static models due to their computational complexity, such that the paper is based on an aggregation of the German electricity network into 28 nodes. However, the approach complements static models taking into account the long-run effects of prices on investment. Several articles discuss the impact of price splits on transmission and generation investment incentives and emission reductions (Grimm et al.,, 2016, 2021, 2022). For example, Grimm et al., (2021) predict that a split of the German market into two price zones has only a small impact on efficiency.
Another stream in the literature asks how zones should be split. While some papers focus on identifying configurations by clustering power transfer distribution factors (PTDFs) (e.g., (Kumar et al.,, 2004) and (Kłos et al.,, 2014)), most works use LMPs as clustering features (Stoft,, 1997). Burstedde, (2012) computes LMPs for the scenario years 2015 and 2020 considering a simplified 72-node representation of the European grid, and evaluates the shapes and sizes of the zones obtained by clustering these LMPs using hierarchical clustering. Breuer et al., (2013) implement a genetic algorithm and apply it on a dataset consisting of 380 kV nodes considering a projected system for 2016 and 2018. Based on six scenarios, Felling & Weber, (2018) compute robust price zone configurations for Central Western Europe by employing hierarchical clustering based on the LMPs of approximately 2,200 nodes and assigning weights to nodes depending on their demand- and supply situation. In a complementary paper, Dobos et al., (2024) use the BZR data and the published nodal prices to compute clusters with low price variance. They show that the clusterings they found are not robust across time or the clustering method used.
Finally, there is an extensive literature on nodal pricing and different pricing rules to deal with the non-convexities of power markets. We refer the interested reader to Liberopoulos & Andrianesis, (2016) for an excellent overview. The most prominent examples are Integer Programming (IP) pricing (O’Neill et al.,, 2005) and Convex Hull (CH) pricing (Gribik et al.,, 2007; Hogan & Ring,, 2003; Bichler et al.,, 2022). IP pricing is used in a number of U.S. markets and so are versions of CH pricing. Both pricing rules require side-payments by the market operator to compensate losses of market participants (make-whole payments, MWPs) or even to compensate all incentives to deviate from the efficient allocation (lost opportunity costs, LOCs). These pricing rules are now being discussed in Europe in the context of non-uniform pricing rules. CH pricing is known to provide poor congestion signals. We also draw on a recent proposal by Ahunbay et al., (2024) that trades off make-whole payments and the quality of the congestion signal and evaluate these three pricing rules in the context of our study, which provides input for the discussion on non-uniform pricing in the EU day-ahead market (All NEMO Committee,, 2023).
@@ 1.3 Contributions
Our goal is to understand how prices in these configurations compare to nodal prices and the current market design with a uniform national price. Specifically, we compare the total costs and prices that arise for national, zonal, and different nodal models. For this we leverage the BZR data set, which provides the most comprehensive and most recent data set available about the European power grid and serves as a foundation for decisions by EU authorities.
In contrast to earlier studies, the average price differences across zonal configurations between 1 and 4 zones for Germany are low, less than 3.5 EUR/MWh. There are also individual days where the price difference between zones is more than 10 EUR/MWh, but on average the differences are small. For the scenarios in the BZR data with climate year 2009, the excess supply in the north of Germany after computing a nodal dispatch was typically such that the cross-zonal lines or interconnectors were not a binding contraint. There was congestion in the network, but not necessarily on the interconnectors. [^5] In 2018, the Austria was split from the joint price zone with Germany. Demand and supply is significantly different between both countries, yet prices (after the gas crises in 2021 and 22) are often remarkably close as well (see https://www.energy-charts.info/). This finding extends to differences in the redispatch between different zonal configurations. It is difficult to model real-world redispatch practices of TSOs. We could only implement two simplified minimal-cost and minimal-volume redispatch methods. Real-world practices are more complicated, but we argue that it provides a reasonable lower bound for redispatch costs. While these redispatch costs decrease slightly from one to two zones, we see little effect with further splits. The cross-zonal capacities play a role for these differences and we will discuss them in more detail below. Our findings do not mean that there cannot be scenarios of supply and demand in Germany where cross-zonal links become binding and there are differences in prices. However, in the BZR dataset, intended to inform ACER and other stakeholders, this is not what we find.
The BZR dataset also provides us with an opportunity for an apples-to-apples comparison of zonal and nodal pricing rules. So, we can quantify the effect of moving from a zonal to a nodal pricing regime. In our study, nodal prices lead to higher generation costs than zonal pricing, but to lower total costs taking our estimated redispatch costs into account. This is due to the fact that redispatch is largely avoided in nodal pricing systems. However, there are different ways how nodal prices can be computed and these nodal pricing rules are currently being discussed in the context of the new capacity management rules CACM 2.0 (All NEMO Committee,, 2023). In other words, the comparison between nodal and zonal prices also depends on the nodal pricing rule. We consider Integer Programming (IP) pricing (O’Neill et al.,, 2005) (as used in U.S. power markets) and Convex Hull (CH) pricing (Hogan & Ring,, 2003; Gribik et al.,, 2007) as two established pricing rules in U.S. markets, and the recently proposed Join pricing rule (Ahunbay et al.,, 2024). The results illustrate the trade-offs that can be expected in terms of quality of the congestion signals and minimum make-whole payments. [^6] We compared the nodal prices that we computed to those published in the BZR dataset. The latter were based on duals of a linear programming relaxation of entire European dispatch problem. Details of this model have not been released but the average nodal prices that we computed are not significantly different to those pubished in the BZR dataset in a sample of hourly prices analyzed. While our analysis focuses on Germany, the policy implications of our study go beyond in that they analyze the differences between different zonal and nodal pricing rules based on a realistic dataset.
1.4 Limitations
@@
Like any empirical analysis of its kind, also our study makes assumptions and faces some limitations. First, a central assumption that all such studies need to make for a split of the price zone are the cross-zonal line capacities. Transmission capacity allocation in the field is a complex task with several decision made by TSOs, that cannot easily be predicted and modeled (see Section 2.2.1). In line with earlier studies (see Section 2.2.1 for a discussion), we compute market clearing solutions in which only 80% of the transmission line limits are considered to take into account security considerations. With higher security margins there might be higher differences in the prices between zones. Second, as in any simulation, the way how we compute redispatch costs can only be an approximation of real-world practices, which are complex (see decision BK8-18-0007-A by the German Bundesnetzagentur). For example, we do not provide a detailed accounting of curtailment costs. We also assume that every generator can be redispatched, but this is not the case in reality. So, the computed redispatch costs can only provide a lower bound, but we argue that it provides a fair comparison. Third, our examination focuses on the German day-ahead market without incorporating neighboring countries, cross-border trades, or loop flows through other countries. We do incorporate 100 neighbouring nodes not in Germany though to mitigate the effect, and argue that the consideration of more nodes outside Germany will not lead to higher differences in the prices between zones in Germany but rather the opposite. Fourth, we do not consider investment incentives and the long-run effects of price differences. Such questions are considered in multi-level game-theoretical models, but these incentives also contribute to more similar prices rather than price differences across the zones (Ambrosius et al.,, 2020). Fifth, we assume fixed demand from the ENTSO-E dataset, neglecting the potential for demand response under nodal prices. Incentives for demand response would be in favour of nodal pricing rules. Sixth, the current European market is characterized by portfolio bidding, differing from the unit-commitment bids underlying both the study by ENTSO-E and our analysis. Portfolio bids allow optimization of trades across assets and differ from those feasible in nodal market designs. Incorporating such considerations into our analysis would require making strong and potentially unwarranted assumptions regarding market participant behavior. Neither the LMP study nor we make such assumptions. Seventh, as any other study of this sort, we do not model the intraday or forward markets. Finally, the data set provided by ENTSO-E was constructed for the target year 2025, and it is important to acknowledge that analyses can vary based on differing assumptions regarding supply and demand. However, the ENTSO-E dataset represents the best source available today for such a study. The assumptions made for our analysis are described in Section 3.
We also want to emphasize that the discussion surrounding zonal and nodal pricing extends beyond dispatch and prices. In the ongoing Bidding Zone Review, a comprehensive approach is adopted which considers factors such as liquidity, security of supply, transition costs, and investments in low-carbon technologies, which we can not replicate in our analysis.
1.5 Organization
@@ 2.2 Market Clearing
In practice, market operators employ different bidding languages to encode $v_{b}$ and $c_{s}$, as well as different approximations $\Psi$ of the power flow constraints. In this work, we use the data released for the European bidding zone review to compare allocation and prices under different market clearing mechanisms.
In full generality, the market operator seeks an allocation $(x,y)$, where $x=(x_{b})_{b\in B}$, $x_{b}\in\mathbb{R}^{N\times T}$ is the allocation vector of buyers and $y=(y_{s})_{s\in S}$,$y_{s}\in\mathbb{R}^{N\times T}$ is the allocation vector of sellers. The market operator identifies an allocation that maximizes welfare, by solving the following optimization model:
@@
On European day-ahead markets, buyers and sellers submit hourly bids and block orders (NEMO Committee,, 2019). The BZR data, however, does not contain bids of this structure. Instead, it provides generator and load characteristics similar to unit commitment problems in U.S. markets. Specifically, for a buyer $b$ located in a node $n_{b}$ [^7] For $t\in T,n\in N\setminus\{n_{b}\},x_{b,nt}=0$. there is only information on a fixed demand profile $P_{b}\in\mathbb{R}^{N\times T}$, while price-elastic bids are unavailable. Consequently, the valuation function of $b$ simplifies to strict demand satisfaction.
@@
Given that only price-inelastic demand bids are available, the objective of problem (1) can be interpreted as minimizing generators’ costs required to fulfill the demand.
@@
$\Psi^{DC}=\{(x,y):\sum_{s\in S}y_{s,nt}-\sum_{b\in B}x_{b,nt}=\sum_{m\in N}-B_{nm}(\theta_{nt}-\theta_{mt})\;\forall n\in N,t\in T\}$ (4)
In (4), $B_{nm}$ is the susceptance of the line connecting nodes $n$ and $m$. The voltage angle decision variable for node $n$ and period $t$ is denoted as $\theta_{nt}$. If such a line does not exist, then $B_{nm}=0$, and $B_{nn}=0\;\forall n$. The constraint set in (4) can be enriched by further constraints, such as box constraints on the voltage angles, and more (Li et al.,, 2022). We use such a standard DCOPF model in our analysis, similar to what was used in the ENTSO-E study. Note that this representation of the DCOPF using line susceptances is equivalent to the representation that considers PTDFs. When the representation $\Psi$ of the transmission network includes all nodes in the transmission network, we refer to nodal market clearing. When only a subset of the grid is considered, we speak of zonal market clearing, as discussed in the following. Power system operators need to consider contingencies for a secure operation of the network. The system should satisfy the ($N-1$) criterion to guarantee that the power system is operated safely in the normal condition or in any contigency case (Christie et al.,, 2000; Marien et al.,, 2013; Weinhold & Mieth,, 2024). Rather than a full security-constrained optimal power flow (SCOPF) model (Hinojosa,, 2020), we consider a safety margin of 20% on all transmission lines.